# Can Large Language Models Faithfully Express Their Intrinsic Uncertainty in Words?

### v1

@@ 1 Introduction
However, uncertainty expressions are only useful when they faithfully reflect the model’s intrinsic uncertainty. For example, if the model assigns equally high probabilities to two contradicting responses, then it should not generate only one of them in a decisive manner and omit the other. We formalize this through the notion of faithful response uncertainty2), an example-level score that quantifies the gap between the (linguistic) decisiveness in which the model conveys its assertions and its intrinsic confidence in them ( Fig. 1). Our approach differs significantly from prior work in that we aim to align the decisiveness of the generated response with the model’s intrinsic confidence, rather than with external factuality verdicts Kadavath et al. (2022); Kuhn et al. (2023); Lin et al. (2022); Mielke et al. (2022); see discussion in §C and additional related work in §D.
@@ 3 Measuring Decisiveness & Uncertainty
Following previous work Kuhn et al. (2023); Manakul et al. (2023); Tian et al. (2023a), we quantify certainty via consistency. Concretely, for a query $\mathbf{Q}$ (e.g., “When was Barack Obama born?”) we quantify the uncertainty of a generated assertion $A$ (e.g., “Barack Obama was born in 1961”) by examining the consistency between this assertion and re-sampled answers to $\mathbf{Q}$: If the generated answers agree with $A$ (e.g., “1961”, “I think he was born in 1961”, or “August 4, 1961.”), then we say $M$ is confident in $A$. Conversely, assertions that contradict $A$ (e.g., “1962” or “Probably 1955”) indicate that $M$’s confidence in $A$ is lower. [^2] Notably, this formulation modestly deviates from Kuhn et al. (2023), which we explain in §C. Formally, given a question $\mathbf{Q}$ and a generated response $\mathbf{R}$ consisting of a single assertion $A$, let $\left\{\mathbf{R}_{1},\dots,\mathbf{R}_{k}\right\}$ be the set of sampled responses and $\left\{A_{1},\dots,A_{k}\right\}$ the set of corresponding assertions (i.e., $\mathcal{A}(\mathbf{R}_{i})=\{A_{i}\}$). We quantify the confidence of $M$ in $A$ as the fraction of sampled assertions that contradict $A$:
@@
We implement the above scores (Eq. 1, Eq. 2) by prompting a “judge” LLM. For a given query $\mathbf{Q}$ and a generated response $\mathbf{R}$, we first extract the assertion $A$ in $\mathbf{R}$ and its decisiveness score using a few-shot prompt $\mathcal{P}_{d}$ (see Tab. 5 in §E). Next, to quantify the model’s intrinsic confidence, we sample $k=20$ additional answers for $\mathbf{Q}$ and extract their corresponding assertions with $\mathcal{P}_{d}$. Then, we use another few-shot prompt $\mathcal{P}_{c}$ (see Tab. 6 in §E) to check for every extracted assertion whether it contradicts $A$. In our experiments, we use Gemini Ultra as the judge.
@@ 4 Experimental Setting
As we focus on model uncertainty, [^3] Our formulation also extends to data uncertainty, see §C. we exclude ambiguous questions for which uncertainty can rise due to data uncertainty (see details in §A).
@@ Methods
We obtain answers using multiple prompts (see §E) with greedy decoding:
@@
Given a model $M$ and a set of QA pairs $\left\{(\mathbf{Q}_{i},\mathbf{R}_{i})\right\}_{i=1}^{n}$, the mean faithful generation metric ($\mathtt{MFG}$) quantifies the expected faithfulness of a single answer: $\mathtt{MFG}=\frac{1}{n}\sum_{i=1}^{n}\left[\mathtt{f}_{M}(\mathbf{R}_{i};\mathbf{Q}_{i})\right]$. While $\mathtt{MFG}$ is a useful indicator, it heavily depends on the distribution of confidence values the model admits [^4] E.g., if $\mathtt{dec}\equiv 1$, $\mathtt{MFG}$ will be the mean confidence value., making it less useful for comparing different models. Therefore, we utilize a second metric, conditional mean faithful generation ($\mathtt{cMFG}$), that additionally conditions on the confidence level: $\textbf{E}_{\begin{subarray}{c}{i\sim n}\\ {v\sim U[0,1]}\end{subarray}}\left[\mathtt{f}_{M}(\mathbf{R}_{i};\mathbf{Q}_{i})\,|\,\mathtt{conf}_{M}(\mathbf{R}_{i};\mathbf{Q}_{i})=v\right]$. In practice, we bin the conf. scores to 10 equally-sized bins and condition on each bin. Note that $\mathtt{cMFG}$ essentially simulates $\mathtt{MFG}$ with uniformly random confidence scores, making it more appropriate for comparing different models. Particularly, $0.5$ is a baseline value for $\mathtt{cMFG}$ as it is obtained for two simple decisiveness strategies that are independent of the model’s confidence (always answering decisively / at a random level of decisiveness). We additionally track mean accuracy, decisiveness, confidence, and punting rate.
@@ 5 Results
Without special instructions, models generate decisive answers, even for uncertain answers (Fig. 3).
Considering the Vanilla baseline, all models perform poorly in terms of faithfulness with $\mathtt{cMFG}$ close to 0.5 (Tab. 1, top row). Specifically, models do not generate any expressions of uncertainty (i.e., $\mathtt{dec}=1$), irrespective of the model’s intrinsic confidence in the answer.
@@
For the non-Vanilla baselines, there is a small increase in $\mathtt{cMFG}$, with maximal scores reaching $0.63$ for GPT-4 and $0.7$ and Gemini-Ultra. While prompting models to express uncertainty sometimes yields hedging expressions (see examples in Tab. 3 in §B), the correlation between decisiveness and confidence is weak, indicating LLMs often hedge confident answers and answer decisively despite uncertainty (see Fig. 4 in §B).
Figure 3: Mean accuracy, confidence, and decisiveness scores for Vanilla on PopQA (results on NQ show similar trends, see §B). Even the most accurate models answer decisively, despite non-trivial uncertainty.
@@ 6 Conclusion
We formalize the desiderata that a model’s generated response should reflect its intrinsic uncertainty in natural language. Our evaluations reveal that modern LLMs perform poorly at this task, stressing the need for better alignment techniques towards ensuring trustworthiness in LLMs.
@@ Appendix A Implementation details
As mentioned in §3, we implement our decisiveness score (Eqn. 1) with a few-shot prompt $\mathcal{P}_{d}$ in which the LLM is instructed to extract the assertion and a score between $0.0$ and $1.0$ (see Tab. 5 in §E), judging purely based off of the provided response. This is aligned with the output of the confidence score (which is naturally a number on $[0,1$), and gives rise to a numeric faithfulness score. We also explored obtaining the numeric decisiveness score by scoring the probability of a positive answer to the question Is this statement decisive? (similar to P(True) in Kadavath et al. (2022)), but this did not work as well, and also has the disadvantage of requiring access to the probability of generated tokens, which is not always available in public-facing APIs.
@@
The prompts we use for decisiveness and confidence scoring can be found in Tab. 5 and 6, respectively. In Tab. 2 we provide the specific instructions in the methods we test.
Method Instruction
@@ Appendix B Results
Figure 6: Standard decoding yields decisive answers, even under uncertainty: We show Accuracy (blue), Confidence (orange) and Decisiveness (green) for each model (x-axis) and dataset, for standard decoding. The x-axis is sorted according to Accuracy. We see: (1) More accurate models generally tend to have higher confidence. (2) Even the best models have some significant uncertainty (e.g. on the challenging PopQA benchmark, the high confidence is  0.8). (3) All the models answer decisively, regardless of their uncertainty.
In Fig. 6 we plot the mean accuracy, confidence and decisiveness scores for each model-dataset combination, for regular decoding (Vanilla). In Fig. 5 we show the mean confidence vs decisiveness for the other decoding methods. In Tab. 3 we show random examples in which the uncertainty prompt induces answers with hedging expressions. In Fig. 4 we plot decisiveness vs confidence for two of the best performing model-method-dataset combinations, revealing the correlation between confidence and decisiveness is weak.
Appendix C Discussion
@@
There are different sources of uncertainty: epistemic (“data uncertainty”, e.g. when the user’s intent or the question is ambiguous) and aleatoric (“model uncertainty”, where the model itself may lack perfect knowledge to answer the question). The source of the uncertainty can be important in determining the optimal approach for conveying the uncertainty linguistically Juanchich et al. (2017); see Table 4 for examples. While our framework and definitions are generally applicable, in this work our focus is on model uncertainty.
@@
Appendix D Related work
@@
There are many different approaches for eliciting factuality-based confidence scores from LLMs. Unsupervised approaches involve examining the agreement across multiple answers Kuhn et al. (2023); Manakul et al. (2023); Tian et al. (2023a), probing the internal representations Azaria and Mitchell (2023); Burns et al. (2022) or directly prompting the model Kadavath et al. (2022). This line of work differs from our work in that we focus on expressing uncertainty linguistically rather than eliciting it post-generation, and that our evaluation hinges on faithfulness rather than factuality (see elaborated discussed in §C).
@@
Lin et al. (2022) show that GPT3.5 can be trained to output calibrated “verbalized confidence” (i.e., generate its level of confidence in language, e.g. “61%” or “medium confidence”) [^5] Specifically, they use a weak signal for the predicted accuracy of an answer - the mean training accuracy of the model at the task for examples of that “type” (e.g., 2 digit numbers). on a suite of arithmetic tasks. In a similar vein, Tian et al. (2023b) show that verbalized confidence outperforms token-based confidence elicitation when the evaluated LMs are fine-tuned with reinforcement learning from human feedback. Our faithfulness objective is different from “verbalized uncertainty” in that it requires the model to naturally incorporate the uncertainty into the response itself, which is significantly more expressive (and more closely resembles how humans communicate uncertainty). Finally, Mielke et al. (2022) is closely related to our work: They propose a controllable-generation framework, in which the dialogue agent’s response is adjusted by choosing linguistic control tokens based on the predicted probability that the chatbot’s answer is correct. Beyond the differences discussed in §C, our approach is also different in two additional aspects. First, our decisiveness notion is numeric and can therefore capture distinctions within the vast landscape of hedging expressions. For example, both “I’m almost sure..” and “I have no idea, …” express some uncertainty and hence their binary approach groups these examples together, whereas our decisiveness scoring will assign them very different scores. Second, their evaluation is on older LMs (small-scale and non-instruction-tuned), that have very low accuracy to begin with (e.g. 5% on TriviaQA).
@@
Appendix E Prompts used in Experiments

### v2

@@ Abstract
Can Large Language Models Faithfully Express
Their Intrinsic Uncertainty in Words?
Gal Yona Google Research galyona@google.com          Roee Aharoni Google Research roeeaharoni@google.com          Mor Geva Tel Aviv University, Google Research pipek@google.com
@@ 1 Introduction
However, uncertainty expressions are only useful when they faithfully reflect the model’s intrinsic uncertainty. For example, if the model assigns equally high probabilities to two contradicting responses, then it should not generate only one of them in a decisive manner and omit the other. We formalize this through the notion of faithful response uncertainty2), an example-level score that quantifies the gap between the (linguistic) decisiveness in which the model conveys its assertions and its intrinsic confidence in them ( Fig. 1). Our approach differs significantly from prior work in that we aim to align the decisiveness of the generated response with the model’s intrinsic confidence, rather than with external factuality verdicts Kadavath et al. (2022); Kuhn et al. (2023); Lin et al. (2022); Mielke et al. (2022); see discussion in §B and additional related work in §C.
@@ 3 Measuring Decisiveness & Uncertainty
Following previous work Kuhn et al. (2023); Manakul et al. (2023); Tian et al. (2023a), we quantify certainty via consistency. Concretely, for a query $\mathbf{Q}$ (e.g., “When was Barack Obama born?”) we quantify the uncertainty of a generated assertion $A$ (e.g., “Barack Obama was born in 1961”) by examining the consistency between this assertion and re-sampled answers to $\mathbf{Q}$: If the generated answers agree with $A$ (e.g., “1961”, “I think he was born in 1961”, or “August 4, 1961.”), then we say $M$ is confident in $A$. Conversely, assertions that contradict $A$ (e.g., “1962” or “Probably 1955”) indicate that $M$’s confidence in $A$ is lower. [^2] Notably, this formulation modestly deviates from Kuhn et al. (2023), which we explain in §B. Formally, given a question $\mathbf{Q}$ and a generated response $\mathbf{R}$ consisting of a single assertion $A$, let $\left\{\mathbf{R}_{1},\dots,\mathbf{R}_{k}\right\}$ be the set of sampled responses and $\left\{A_{1},\dots,A_{k}\right\}$ the set of corresponding assertions (i.e., $\mathcal{A}(\mathbf{R}_{i})=\{A_{i}\}$). We quantify the confidence of $M$ in $A$ as the fraction of sampled assertions that contradict $A$:
@@
We implement the above scores (Eq. 1, Eq. 2) by prompting a “judge” LLM. For a given query $\mathbf{Q}$ and a generated response $\mathbf{R}$, we first extract the assertion $A$ in $\mathbf{R}$ and its decisiveness score using a few-shot prompt $\mathcal{P}_{d}$ (see Tab. 5 in §D). Next, to quantify the model’s intrinsic confidence, we sample $k=20$ additional answers for $\mathbf{Q}$ and extract their corresponding assertions with $\mathcal{P}_{d}$. Then, we use another few-shot prompt $\mathcal{P}_{c}$ (see Tab. 6 in §D) to check for every extracted assertion whether it contradicts $A$. In our experiments, we use Gemini Ultra as the judge.
@@ 4 Experimental Setting
As we focus on model uncertainty, [^3] Our formulation also extends to data uncertainty, see §B. we exclude ambiguous questions for which uncertainty can rise due to data uncertainty (see details in §A).
@@ Methods
We obtain answers using multiple prompts (see Tab. 2 for the specific instructions) with greedy decoding:
@@
Given a model $M$ and a set of QA pairs $\left\{(\mathbf{Q}_{i},\mathbf{R}_{i})\right\}_{i=1}^{n}$, the mean faithful generation metric ($\mathtt{MFG}$) quantifies the expected faithfulness of a single answer: $\mathtt{MFG}=\frac{1}{n}\sum_{i=1}^{n}\left[\mathtt{f}_{M}(\mathbf{R}_{i};\mathbf{Q}_{i})\right]$. While $\mathtt{MFG}$ is a useful indicator, it heavily depends on the distribution of confidence values the model admits [^4] E.g., if $\mathtt{dec}\equiv 1$, $\mathtt{MFG}$ will be the mean confidence value., making it less useful for comparing different models. Therefore, we utilize a second metric, conditional mean faithful generation ($\mathtt{cMFG}$), that additionally conditions on the confidence level: $\textbf{E}_{\begin{subarray}{c}{i\sim n}\\ {v\sim U[0,1]}\end{subarray}}\left[\mathtt{f}_{M}(\mathbf{R}_{i};\mathbf{Q}_{i})\,|\,\mathtt{conf}_{M}(\mathbf{R}_{i};\mathbf{Q}_{i})=v\right]$. In practice, we bin the conf. scores to 10 equally-sized bins and condition on each bin. Note that $\mathtt{cMFG}$ essentially simulates $\mathtt{MFG}$ with uniformly random confidence scores, making it more appropriate for comparing different models. Particularly, $0.5$ is a baseline value for $\mathtt{cMFG}$ as it is obtained for two simple decisiveness strategies that are independent of the model’s confidence (always answering decisively / at a random level of decisiveness).
Figure 3: Standard decoding yields decisive answers, even under uncertainty: We show results for standard decoding on PopQA (left) and NQ (right). Models (x-axis) are sorted by Accuracy (blue), and the additional bars show Confidence (orange) and Decisiveness (green). We see: (1) More accurate models generally tend to have higher confidence. (2) Even the best models have some significant uncertainty (e.g. on the challenging PopQA benchmark, the high confidence is  0.8). (3) All the models answer decisively, regardless of their uncertainty.
@@ 5 Results
Without special instructions, models generate decisive answers, even for uncertain answers
Considering the Vanilla baseline, all models perform poorly in terms of faithfulness with $\mathtt{cMFG}$ close to 0.5 (Tab. 1, top row). This happens because all the models we tested did not generate any expressions of uncertainty, despite having significant intrinsic uncertainty in some cases. See Fig. 3.
@@
For the non-Vanilla baselines, there is a small increase in $\mathtt{cMFG}$, with maximal scores reaching $0.63$ for GPT-4 and $0.7$ and Gemini-Ultra. We observe that prompting models to express uncertainty slightly reduces the mean decisiveness (Fig. 5) by introducing hedging expressions (see examples in Tab. 3 in the Appendix). Importantly, however, the correlation between decisiveness and confidence is weak; see Fig. 4. This suggests that LLMs hedge when they are confident and answer decisively despite uncertainty, explaining the still-low $\mathtt{cMFG}$ scores.
@@ 6 Conclusion
We formalize the desiderata that a language model’s generated response should reflect its intrinsic uncertainty in natural language. We instantiate this with a generic evaluation framework that quantifies the mismatch between the linguistic decisiveness of the assertions made in the response and the intrinsic uncertainty the model has in these assertions. We quantify intrinsic uncertainty by considering the consistency of model answers across multiple samples, but our framework is generic and can be extended to other methods for quantifying uncertainty, such as methods that depend on the internal representations of the model.
We evaluate the abilities of modern LLMs at the task of faithfully conveying their intrinsic uncertainty. Taken together, our evaluations reveal that modern LLMs perform poorly at this task, stressing the need for better alignment techniques towards ensuring trustworthiness in LLMs.
Limitations
Determining the “appropriate” way for language models to convey uncertainty in natural language is a complex and multi-faceted question, with deep roots in computational linguistics (see §C). To simplify our evaluation, we focused on one relatively simple case for conveying uncertainty: explicitly conveying aleatoric uncertainty (or “model uncertainty”) in knowledge-intensive QA tasks using hedging expressions. Technically, this was facilitated by evaluating on non-ambiguous questions (namely, questions that have a single correct answer). For Natural Questions Kwiatkowski et al. (2019) we relied on prior work by Min et al. (2020) that identified its non-ambiguous subset, but for PopQA Mallen et al. (2022) we relied on heuristics (such as choosing a subset of relations and keeping only questions with a single correct answer in the dataset), which may not be perfect. On a more conceptual level, the fact that we show overwhelmingly negative results even for this simple setting serves to highlight what is, in our opinion, a significant limitation of modern LMs. As LMs improve at faithfully conveying uncertainty is simple settings, exploring the intricacies involved with expressing uncertainty in more generic settings (data uncertainty, implicitly vs explicitly, etc) will become more important.
We conclude with a discussion of two aspects regarding our evaluations, in which we demonstrated that the ability to steer modern language models towards faithfully conveying their uncertainty via prompting is limited. First, since we show an overwhelmingly negative result regarding the limitation of existing LLMs, we chose to focus on the best-performing model families (Gemini Gemini-Team (2023) and GPT Achiam et al. (2023)). As such, our evaluation uses closed models (i.e., models that are available via a public API, but whose weights are proprietary). Second, the methods we tested relied on zero-shot and few-shot prompts. Concurrently to our work, Agarwal et al. (2024); Bertsch et al. (2024) introduced the many-shot regime for prompting models with extremely long context windows, showing promising results. Understanding whether this creates a fundamental difference in the ability of language models to faithfully convey uncertainty is another interesting direction for future work.
@@ Appendix A Implementation details
As mentioned in §3, we implement our decisiveness score (Eqn. 1) with a few-shot prompt $\mathcal{P}_{d}$ in which the LLM is instructed to extract the assertion and a score between $0.0$ and $1.0$ (see Tab. 5 in §D), judging purely based off of the provided response. This is aligned with the output of the confidence score (which is naturally a number on $[0,1$), and gives rise to a numeric faithfulness score. We also explored obtaining the numeric decisiveness score by scoring the probability of a positive answer to the question Is this statement decisive? (similar to P(True) in Kadavath et al. (2022)), but this did not work as well, and also has the disadvantage of requiring access to the probability of generated tokens, which is not always available in public-facing APIs.
@@
The prompts we use for decisiveness and confidence scoring can be found in Tab. 5 and 6, respectively.
@@
Appendix B Discussion
@@
There are different sources of uncertainty: epistemic (“data uncertainty”, e.g. when the user’s intent or the question is ambiguous) and aleatoric (“model uncertainty”, where the model itself may lack perfect knowledge to answer the question). The source of the uncertainty can be important in determining the optimal approach for conveying the uncertainty linguistically Juanchich et al. (2017) (see Table 4 for examples). While our framework and definitions are generally applicable, in this work our focus is on model uncertainty. Exploring uncertainty communication in the context of aleatoric uncertainty is an interesting direction for future work, especially as prior work suggests noticeable differences between the variability of text continuation between humans and LLMs Giulianelli et al. (2023).
@@
Appendix C Related work
@@
There are many different approaches for eliciting factuality-based confidence scores from LLMs. Unsupervised approaches involve examining the agreement across multiple answers Kuhn et al. (2023); Manakul et al. (2023); Tian et al. (2023a), probing the internal representations Azaria and Mitchell (2023); Burns et al. (2022) or directly prompting the model Kadavath et al. (2022). This line of work differs from our work in that we focus on expressing uncertainty linguistically rather than eliciting it post-generation, and that our evaluation hinges on faithfulness rather than factuality (see elaborated discussed in §B).
@@
Lin et al. (2022) show that GPT3.5 can be trained to output calibrated “verbalized confidence” (i.e., generate its level of confidence in language, e.g. “61%” or “medium confidence”) [^5] Specifically, they use a weak signal for the predicted accuracy of an answer - the mean training accuracy of the model at the task for examples of that “type” (e.g., 2 digit numbers). on a suite of arithmetic tasks. In a similar vein, Tian et al. (2023b) show that verbalized confidence outperforms token-based confidence elicitation when the evaluated LMs are fine-tuned with reinforcement learning from human feedback. Our faithfulness objective is different from “verbalized uncertainty” in that it requires the model to naturally incorporate the uncertainty into the response itself, which is significantly more expressive (and more closely resembles how humans communicate uncertainty). Finally, Mielke et al. (2022) [^6] And Krause et al. (2023) in a multi-lingual setting. is closely related to our work: They propose a controllable-generation framework, in which the dialogue agent’s response is adjusted by choosing linguistic control tokens based on the predicted probability that the chatbot’s answer is correct. Beyond the differences discussed in §B, our approach is also different in two additional aspects. First, our decisiveness notion is numeric and can therefore capture distinctions within the vast landscape of hedging expressions. For example, both “I’m almost sure..” and “I have no idea, …” express some uncertainty and hence their binary approach groups these examples together, whereas our decisiveness scoring will assign them very different scores. Second, their evaluation is on older LMs (small-scale and non-instruction-tuned), that have very low accuracy to begin with (e.g. 5% on TriviaQA).
@@
Appendix D Prompts used in Experiments